A steam turbine develops 149 kW while receiving 930 kJ from each kilogram of steam passing through the turbine. Calculate the kilograms of steam per hour required to produce this power (in kg/hr).
1
Expert's answer
2020-03-23T10:35:28-0400

"P=\frac{Q}{t}=\frac{q\cdot m}{3600\cdot n}\to \frac{m}{n}=\frac{P\cdot3600}{q}=\frac{149000\cdot3600}{930000}\approx576.8" "kg\/hr"